Find the class boundaries midpoints and widths for the given class 43-47


1
Expert's answer
2021-11-09T16:20:07-0500

Class boundaries: 43-0.5=42.5 and 47+0.5=47.5

Midpoint: "\frac{43+47}{2}=45"

Width: 47-43=4